### v2

@@ 1 Introduction
Figure 1: Performance comparison on the MathVerse benchmark. Our AStar framework achieves competitive results against most open-sourced MLLMs and closed-source ones, showing outstanding structured thinking and reasoning abilities.
Multimodal Large Language Models (MLLMs) have demonstrated impressive capabilities across diverse tasks and domains Yin et al. (2023); OpenAI (2024); Chen et al. (2024c), such as autonomous driving Cui et al. (2024), and visual question answering Hartsock & Rasool (2024). Their proficiency in complex multimodal reasoning, particularly in mathematical tasks involving visual content, has emerged as a critical benchmark for evaluating fundamental cognitive abilities towards strong artificial intelligence Searle (1980); Goertzel & Pennachin (2007); Wang et al. (2024f); Qiao et al. (2024). Mastering multi-step visual reasoning requires the integration of multimodal information, along with rigorous adherence to complex rules and sophisticated problem-solving strategies, presenting significant challenges for existing MLLMs Zhang et al. (2024g); Wang et al. (2024a).
Figure 2: Schematic comparison between AStar and two mainstream structured reasoning methods. $\mathcal{L}_{(\cdot)}$ denotes the training optimization objective. (a) Search-based methods suffer from computational inefficiency due to extensive solution space iterations. (b) Teacher-guided training optimizes models using distilled rationales from powerful models like GPT-4o, but requires substantial data and computational resources, resulting in low-efficiency pattern extraction and poor data utilization. (c) Our approach effectively combines MLLMs’ internal implicit reasoning capabilities with explicitly extracted insights, achieving a compelling balance between performance and efficiency.
Inspired by recent advances in System 2 slow-thinking reasoning systems like OpenAI o1 OpenAI (2024) and QVQ Team (2024), there is growing interest in incorporating structured thinking into MLLMs Xu et al. (2024); Dong et al. (2024d); Du et al. (2025). This direction aims to address the limitations of conventional MLLMs that often rely on simple ‘direct prediction’ modes due to the scarcity of high-quality long-chain reasoning data Xu et al. (2024); Luo et al. (2025). According to existing literature, two primary approaches have emerged for implementing slow-thinking reasoning systems: explicit search and teacher supervision-guided training. The first approach leverages explicit search structures (e.g., Monte Carlo tree search, MCTS) with specialized reward models to guide the exploration of solution paths Dong et al. (2024a); Yao et al. (2024). The second approach focuses on distilling structured reasoning patterns through long-form Chain-of-Thought (CoT) Wei et al. (2022); Zhang et al. (2024g); Luo et al. (2025) instruction data, typically requiring supervision from closed-source models like GPT-4o for data synthesis.
Despite these advances, current reasoning paradigms face three critical limitations (Figure 2). First, search-based methods Dong et al. (2024a) suffer from computational inefficiency due to extensive solution space iterations. Second, teacher-guided training methods Hu et al. (2024); Yao et al. (2024); Luo et al. (2025) typically require substantial training data ($\geq$100K) and computational resources to implicitly extract reasoning patterns, resulting in low efficiency and poor data utilization. They also heavily depend on proprietary models like GPT-4o for data synthesis, making them impractical for researchers outside major enterprises. Third, static and predefined reasoning processes Xu et al. (2024); Thawakar et al. (2025) constrain flexibility, leaving the reasoning potential of MLLMs underexplored.
To address these challenges, we propose AStar, an Automated Structured Thinking paradigm for multimodAl Reasoning via MCTS. Our approach introduces a novel mechanism for automatically deriving high-level cognitive reasoning patterns from limited data (500 samples) using MCTS-powered hierarchical structures (aiming to address the second limitation). Building on these explicit patterns, we design a unified reasoning framework that seamlessly integrates internal and external advantageous attributes, enabling efficient adaptive inference with minimal tree iterations, thereby tackling the first and third limitations. This novel reasoning paradigm effectively combines MLLMs’ internal implicit reasoning capabilities with external explicit reasoning guidelines, achieving a compelling balance between performance and efficiency.
Specifically, our method comprises three steps: (1) visual reasoning action definition, (2) MCTS-powered thought card construction, and (3) adaptive reasoning and verification. First, we define six atomic reasoning actions as building blocks of chain-structured reasoning patterns (termed “thought cards” that serve as reference insights during inference). These actions simulate human-like cognitive behaviors, including problem decomposition and reasoning step reflection. Second, using a small seed dataset (500 samples), we apply MCTS to derive reference reasoning patterns to construct multiple thought cards. Finally, in the reasoning stage, we select the five optimal thought cards most aligned with the target problem’s cognitive complexity. With these reasoning guidelines, we perform visual reasoning and validate the final solutions via self-consistency checks or outcome reward models. Experiments demonstrate that AStar exhibits impressive reasoning performance with enhanced efficiency, comparable to powerful closed-source models like GPT-4o (Figure 1). Our main contributions are:
Figure 3: Flowchart of our proposed method AStar. This framework consists of three main parts: (1) Visual Atomic Reasoning Action Definition; (2) MCTS-Powered Thought Card Construction; (3) Adaptive Reasoning and Verification.
  • Automated Reasoning Paradigm: proposing an MCTS-based automated approach for generating and selecting the optimal reasoning patterns.
  • Efficient Action-Chain Guided Reasoning: providing explicit guidance for each step of the visual reasoning process, enhancing structured thinking capabilities.
@@
  • Improved Efficiency: achieving comparable performance to recent tree-based methods while reducing inference overhead by 6.4×6.4$\times$6.4 ×, and matching training-based methods while requiring 520×520$\times$520 × less prior data.
2 Related Work
Multimodal Reasoning   Recent advancements in MLLMs have demonstrated robust capabilities across diverse domains, including visual understanding Zhang et al. (2024c), mathematics Zhang et al. (2024d); Du et al. (2025), and scientific inquiries Zong & Qiu (2024). Despite these achievements, complex multimodal reasoning remains challenging due to its demands on both visual perception and high-level cognition. Inspired by OpenAI o1’s impressive performance, recent approaches Zhang et al. (2024e); Xu et al. (2024); Thawakar et al. (2025) attempt structured reasoning with pre-defined stages, enhancing MLLMs’ CoT capabilities Zhang et al. (2024g). However, their rigid structure limits flexibility across different tasks, overlooking the importance of adaptive reasoning in unleashing multimodal reasoning potential Wang et al. (2024e). Our approach addresses this by introducing a hierarchical tree structure that enables task-specific reasoning path generation and selection.
Tree-based Search   Tree structures have demonstrated significant potential in language models Zhang et al. (2024a); Qi et al. (2024); Wu et al. (2024a). Recent efforts explore applying these tree search methods to search effective reasoning paths for MLLMs. While AR-MCTS Dong et al. (2024a) enhances multimodal reasoning by integrating MCTS with active retrieval, its extensive iteration requirements and computational overhead limit practical applications. Similarly, Mulberry Yao et al. (2024) leverages tree structures to distill 260K long-chain reasoning data from powerful models like GPT-4o, but requires substantial computational resources and high-capacity teacher models. These methods struggle to achieve an optimal balance between performance and efficiency. To address these limitations, we propose incorporating high-level reasoning abstractions into MCTS, achieving competitive performance with higher efficiency.
@@ 3 Methodology
Overview of AStar This section introduces AStar in detail. As shown in Figure 3 and Algorithm 1, our approach consists of three steps:
@@
  • MCTS-Powered Thought Card Construction: Leverage MCTS to systematically construct thought cards, which serve as reference insights during inference.
  • Adaptive Reasoning and Verification: Dynamically select and execute optimal reasoning patterns based on problem complexity, followed by solution verification.
@@ 3.1 Visual Reasoning Action Definition
Understanding human complex reasoning is crucial for modeling cognitive processes Jaffe et al. (2023). Existing studies distinguish between two cognitive systems: System 1 and System 2 Kahneman (2011); Da Silva (2023). While “System 1” represents fast, intuitive, yet error-prone thinking, “System 2” involves slow, deliberative thinking with superior performance. With the emergence of advanced models like OpenAI o1 OpenAI (2024) and QVQ Team (2024), developing efficient “System 2” approaches to emulate human cognitive processes has gained significant research attention Xu et al. (2024); Yao et al. (2024); Thawakar et al. (2025). Inspired by this, we introduce six vision-language reasoning actions to bridge the gap between model reasoning and human cognition: Visual Parsing (VP, $a_{1}$), System Analysis (SA, $a_{2}$), One-Step Thought (OST, $a_{3}$), Chain-of-Thought (CoT, $a_{4}$), Divide and Conquer (DC, $a_{5}$), Self-Reflection (SR, $a_{6}$). Details are provided in Appendix B.1.
3.2 MCTS-Powered Thought Card Construction
Following the action definition, we introduce “thought cards” as structured reasoning templates to guide inference in Sec. 3.3. Using a small seed dataset, we first derive reasoning paths (Phase 1) and then distill them into high-level thought cards (Phase 2). These cards serve as prior insights during inference, providing structured guidance for efficient problem-adaptive reasoning.
Phase 1: Acquire reasoning paths for seed data   As shown in Figure 3, we employ MCTS to iteratively optimize the solution search process, generating high-quality reasoning paths for the seed dataset. This design leverages MCTS’s systematic exploration Ye et al. (2021) and MLLMs’ inherent reasoning capabilities Yin et al. (2023); Wang et al. (2024f). We formulate each multimodal reasoning problem $x$ (consisting of input question and images) as a tree search problem, where $x$ represents the root node and subsequent nodes denote reasoning steps (actions and corresponding outcomes) generated by a policy MLLM $\pi_{\theta}$. We define the state $S_{t-1}$ as the trajectory $x,s_{1},...,s_{t-1}$, where $S_{0}=x$. The next step is sampled as $s_{t}\sim\pi_{\theta}(S_{t-1})$. To guide tree expansion, we define $Q(s)$ as the reward value for node $s$. Initially, all unexplored nodes are assigned $Q(s_{i})=0$. They are updated using a weighted average between the parent’s current value and its child node’s value:
@@ 3.2 MCTS-Powered Thought Card Construction
where $\alpha$ is a discount factor for future rewards. For terminal nodes, following Zhou et al. (2024); Wu et al. (2024a), we adopt the likelihood of self-consistency majority voting as the reward value, enabling supervision-free generalization.
Algorithm 1 Multimodal Reasoning with AStar
  Input: a policy model $\pi_{\theta}$; a multimodal test question $x_{t}$; a set of seed data $D_{s}$
  // 3.1. Visual Reasoning Action Definition
  Initialize action space $A=\left\{a_{1},a_{2},a_{3},a_{4},a_{5},a_{6}\right\}$
  // 3.2. MCTS-Powered Thought Card Construction
  $D\leftarrow\left[\;\right]$; $Cards\leftarrow\left\{\right\}$
  for $(x,y)\in$ $D_{s}$ do
     $\{x,y,P\}\leftarrow\text{MCTS}(\pi_{\theta};x)$
     if found an valid reasoning path then
        Find $p_{\text{best}}$ from $P$
        Add $\{x,p_{\text{best}}\}$ into $D$
        Update $Cards$ from $D$
     end if
  end for
  // 3.3. Adaptive Reasoning and Verification
  $c\leftarrow\text{CardMatch}(Cards;x_{t})$
  $y_{t}\leftarrow\text{ReasonAndVerify}(\pi_{\theta};x_{t};c)$
  Output: the optimal reasoning trajectory $y_{t}$
Specifically, this phase comprises four MCTS operations:
(1) Selection. This operation identifies promising nodes for expansion. Starting from the root node, we iteratively select child nodes using the Upper Confidence Bounds applied to Trees (UCT) Kocsis & Szepesvári (2006) until reaching a leaf node:
$UCT(s)=Q(s)+w\sqrt{\frac{lnN(p)}{N(s)}}$ (2)
@@
(3) Simulation. Starting from the selected node, we iteratively sample and expand nodes until reaching a terminal state (maximum depth or answer node).
@@
Phase 2: Distill paths into thought cards   After executing MCTS, we obtain a tree structure for each seed dataset question, yielding multiple valid reasoning paths that constitute the path set $P$. Inspired by the concept of Value of Computation (VOC) Russell & Wefald (1991), which optimizes the trade-off between computational benefits and costs, we propose a VOC-inspired selection metric to identify the optimal reasoning trajectory from candidate solutions:
@@
where $x$ is the task input, $p_{x}$ represents a candidate reasoning path from $P$, and $k$ balances computational benefits against costs. Here, for simplicity, $R(p_{x}|x)$ denotes the path’s final reward (defined as the leaf node’s Q-value), while $C(p_{x})$ measures the reasoning cost (defined as the number of actions in the sequence).
Then, for each question in seed dataset, we select path $p_{best}$ with highest $Score(x,p_{x})$ to build a Question-path repository $D$ with one-to-one mappings. Inspired by metareasoning principles Russell & Wefald (1991); Roberts & Roberts (2024), which advocate for adaptive reasoning strategies, we distill these question-path pairs into abstract thought cards $\mathbb{C}$ using Problem Condition Complexity (PCC) Lee & Heyworth (2000); Embretson & Daniel (2008), a cognitive complexity metric designed for complex reasoning. Each card represents a high-level reasoning pattern abstracted from multiple prior problems, enabling efficient transfer of reasoning insights during inference in the next step.
@@ 3.3 Adaptive Reasoning and Verification
During inference, given a multimodal test query $x_{t}$, we compute its PCC and perform nearest neighbor matching Muja & Lowe (2014) against pre-constructed thought cards $\mathbb{C}$ to identify the most relevant five cards that best align with its complexity level. The selection process is formalized as:
$NN_{5}(x_{t},\mathbb{C})=\mathop{\arg\min}_{\mathbb{C}_{x_{t}}\subseteq\mathbb{C},|S|=5}{\textstyle\sum_{c\in\mathbb{C}_{x_{t}}}}d(x_{t},c)$ (4)
where $NN_{5}(x_{t},\mathbb{C})\subseteq\mathbb{C}$ is the closest subset to $x_{t}$, determined by the distance $d:\mathbb{R}\times\mathbb{R}\rightarrow\mathbb{R}$. By leveraging these reasoning templates, our method maintains the benefits of tree-structured reasoning without extensive node expansion, enabling efficient generation of candidate solutions.
Identifying the optimal reasoning trajectory among these candidates represents a critical challenge Luo et al. (2025). Given the scarcity of verification models in the visual domain, we employ both self-consistency checks and text-domain outcome reward models to select the final solution.
In summary, our approach can be viewed as an optimized variant of tree search algorithms. Unlike traditional MCTS methods that require extensive exploration, we strategically reduce computational complexity by incorporating prior insights through thought cards. As shown in Figure 3, this enables efficient tree traversal along promising trajectories while maintaining high performance, offering valuable insights for developing efficient reasoning strategies.

### v3

@@ 1 Introduction
Multimodal Large Language Models (MLLMs) have demonstrated impressive capabilities across diverse tasks and domains [1, 2, 3], yet they struggle with complex visual reasoning tasks that require processing multimodal information with sophisticated problem-solving strategies [4, 5]. Recent advances in System 2 slow-thinking reasoning models like OpenAI-o1 [6], DeepSeek-R1 [7], and Kimi-K1.5 [8] have inspired growing interest in incorporating structured long Chain-of-Thought (CoT) [9] thinking into MLLMs [10, 11, 12]. These approaches address limitations of conventional MLLMs that often rely on simple “direct prediction” modes due to the scarcity of high-quality long-chain reasoning data [10, 13]. Current methods can be divided into two primary categories: (i) explicit search methods [14, 15]: leverage algorithms like beam search or Monte Carlo Tree Search (MCTS) with specialized reward models to guide extensive solution path exploration; (ii) post-training methods [1, 16]: develop structured long CoT reasoning capabilities through common techniques such as large-scale Supervised Fine-Tuning (SFT) [4, 13] or Reinforcement Learning with Verifiable Rewards (RLVR) like Group Relative Policy Optimization (GRPO) [17, 18].
However, both kinds of methods face some limitations. (i) Search-based methods [14] suffer from computational inefficiency due to extensive exploration across solution spaces; (ii) SFT-based post-training methods [13, 15, 19] typically require substantial training data ($\geq$100K) and computational resources to implicitly extract reasoning patterns, leading to inefficient data utilization. Their reliance on proprietary models like GPT-4o for data synthesis also creates accessibility barriers for researchers outside major enterprises; (iii) RLVR-based post-training methods [20, 21] encounter challenges with exploration depth and training stability. As shown in [22, 23], these methods primarily function by biasing the model’s output distribution toward reward-maximizing paths without introducing external knowledge. This inherently limits their exploration capacity, resulting in a narrower reasoning capability boundary compared to base models.
Figure 1: Evaluation results on MathVerse. AStar makes 7B models competent problem-solvers, surpassing GPT-4o. Our approach demonstrates consistent effectiveness across multiple model architectures and scales.
To address these limitations, we propose AStar, an Automated Structured thinking paradigm for multimodal reasoning. AStar introduces a novel and versatile reasoning framework that seamlessly combines MLLMs’ internal implicit reasoning capabilities with external explicit reasoning guidelines. Specifically, we design “thought cards”, a general and lightweight library storing abstract high-level thoughts (atomic reasoning action-guided reasoning guidelines), which are abstracted from only 500 prior samples using MCTS and can be generalized across tasks. For each test problem, we adaptively retrieve five optimal thought cards aligned with the problem’s characteristics (complexity and semantics). With these abstract guidelines, we perform visual reasoning and obtain the final solution with unified self-consistency checks and outcome reward scoring.
Our method offers four key advantages: (i) Performance Enhancement: By leveraging high-level thought cards, we adaptively obtain reasoning guidelines based on problem attributes, integrating external guidance with the model’s internal capabilities, thereby enhancing performance (Figure 1); (ii) Efficiency Improvement: Our training-free framework achieves comparable performance to powerful baselines with only 500 prior samples, enhancing both data and computational efficiency; (iii) Scalability and Flexibility: Thought cards can be rapidly constructed for new tasks using minimal task-relevant data; and (iv) Generalization and Plug-and-Play Capability: Beyond multimodal reasoning, our framework benefits general visual perception and understanding tasks. It can also function as a plug-and-play test-time inference method compatible with popular post-training techniques like SFT and GRPO. We summarize our core contributions as follows:
  • We introduce an adaptive training-free reasoning framework that combines MLLMs’ internal implicit capabilities with external explicit guidelines, achieving both effectiveness and efficiency.
  • We propose “thought cards” for storing informative high-level thoughts abstracted from minimal prior samples, which can be adaptively instantiated for specific test problems.
@@
Figure 2: Flowchart of our method AStar. This framework consists of three parts: (1) Visual Reasoning Action Definition; (2) Thought Card Construction; (3) Adaptive Reasoning and Verification.
@@ 2 Methodology
Overview of AStar   This section introduces our proposed method AStar in detail. As shown in Figure 2 and Algorithm 1, our approach consists of three steps:
@@
  • Thought Card Construction: Leverage MCTS to systematically construct thought cards, which serve as reference insights and effectively guide subsequent problem-solving during inference.
  • Adaptive Reasoning and Verification: Dynamically select and execute optimal reasoning patterns based on problem complexity, followed by robust solution verification.
@@ 2.1 Visual Reasoning Action Definition
Understanding human complex reasoning is crucial for modeling cognitive processes [24]. Existing studies distinguish between two cognitive systems: System 1 and System 2 [25, 26]. While “System 1” represents fast, intuitive, yet error-prone thinking, “System 2” involves slow, deliberative thinking with superior performance. With the emergence of advanced models like OpenAI o1 [6] and Kimi-K1.5 [8], developing efficient “System 2” approaches to emulate human cognitive processes has gained significant research attention [15, 27, 28]. Inspired by this, we introduce six vision-language reasoning actions to bridge the gap between model reasoning and human cognition: Visual Parsing (VP, $a_{1}$), System Analysis (SA, $a_{2}$), One-Step Thought (OST, $a_{3}$), Chain-of-Thought (CoT, $a_{4}$), Divide and Conquer (DC, $a_{5}$), Self-Reflection (SR, $a_{6}$). Details are provided in Appendix C.1.
2.2 Thought Card Construction
Based on the above action definitions, we introduce “thought cards” as reasoning templates to guide inference in Section 2.3. Using only 500 prior samples, we first derive their reasoning paths (Phase 1) and then distill them into high-level thought cards (Phase 2). These cards provide structured guidance for efficient problem-adaptive reasoning during inference.
Phase 1: Acquire reasoning paths for seed data   As shown in Figure 2, we employ MCTS to iteratively optimize the solution search process, generating high-quality reasoning paths for the seed dataset. This design leverages MCTS’s systematic exploration and MLLMs’ inherent reasoning capabilities [29, 1]. We formulate each multimodal reasoning problem $x$ (consisting of input question and images) as a tree search problem, where $x$ represents the root node and subsequent nodes denote reasoning steps (actions and corresponding outcomes) generated by a policy MLLM $\pi_{\theta}$. We define the state $S_{t-1}$ as the trajectory $x,s_{1},...,s_{t-1}$, where $S_{0}=x$. The next step is sampled as $s_{t}\sim\pi_{\theta}(S_{t-1})$. To guide tree expansion, we define $Q(s)$ as the reward value for node $s$. Initially, all unexplored nodes are assigned $Q(s_{i})=0$. They are updated using a weighted average between the parent’s current value and its child node’s value:
Algorithm 1 Multimodal Reasoning with AStar
1:a policy model $\pi_{\theta}$; a multimodal test query $x_{t}$; a set of seed data $D_{s}$
2:Initialize action space $A=\{a_{1},a_{2},a_{3},a_{4},a_{5},a_{6}\}$triangleright 2.1. Action Definition
3:Initialize repository $D\leftarrow[\;]$; Cards ${}\mathbb{C}\leftarrow\{\}$
4:for $(x,y)\in D_{s}$ do $\triangleright$ 2.2. Thought Card Construction
5:     Acquire reasoning paths $\{x,P\}\leftarrow\text{MCTS}(\pi_{\theta};x)$
6:     if found at least one valid reasoning path then
7:         Find $p_{\text{best}}$ from $P$ (Equation (3))
8:         Add $\{x,p_{\text{best}}\}$ into $D$
9:         Update $\mathbb{C}$ from $D$
10:     end if
11:end for
12:$\mathbb{C}_{x_{t}}\leftarrow\text{Card\_Match}(\mathbb{C};x_{t})$triangleright 2.3. Adaptive Reasoning and Verification
13:$y_{t}\leftarrow\text{Reason\_And\_Verify}(\pi_{\theta};x_{t};\mathbb{C}_{x_{t}})$
14:the optimal reasoning trajectory $y_{t}$
@@ 2.2 Thought Card Construction
where $\alpha$ is a discount factor for future rewards. For terminal nodes, we adopt the likelihood of self-consistency majority voting as reward value, enabling supervision-free generalization. Specifically, this phase comprises four MCTS operations:
(1) Selection. This operation identifies promising nodes for expansion. Starting from the root node, we iteratively select child nodes using the Upper Confidence Bounds applied to Trees (UCT) [30] until reaching a leaf node:
$UCT(s)=Q(s)+w\sqrt{\frac{\ln N(p)}{N(s)}}$ (2)
@@
(3) Simulation. Starting from the selected node, we iteratively sample and expand nodes until reaching a terminal state (maximum depth or final answer node).
@@
Phase 2: Distill paths into thought cards   After executing MCTS, we obtain a tree structure for each seed dataset question, yielding multiple valid reasoning paths that constitute the path set $P$. Inspired by the concept of Value of Computation (VOC) [31], which optimizes the trade-off between computational benefits and costs, we propose a VOC-inspired selection metric to identify the optimal reasoning trajectory from candidate solutions:
@@
where $x$ is the task input, $p_{x}\in P$ denotes a candidate reasoning path, and $k$ balances benefits against computational costs. Here, $R(p_{x}|x)$ denotes the path’s final reward (defined as the leaf node’s Q-value), while $C(p_{x})$ is the reasoning cost (defined as the number of actions in the sequence).
Then, for each question in the seed dataset, we select the path $p_{best}$ with the highest $Score(x,p_{x})$ to build a Question-path repository $D$ with one-to-one mappings. Inspired by metareasoning principles [31, 32], which advocate for adaptive reasoning strategies, we distill these question-path pairs into abstract thought cards $\mathbb{C}$ as high-level guidelines. Each card is distilled based on two attributes: problem condition complexity (PCC) [33] and text-image semantics (TIS) [14, 34]. PCC represents prior known conditions derived from input image-text pairs $(i,t)$, obtained using a lightweight 2B MLLM. TIS refers to joint semantic representations based on the CLIP model encodings of $(i,t)$:
$E_{x}(i,t)=\frac{E_{I}(i)+E_{T}(t)}{2}$ (4)
Finally, each thought card contains a high-level thought template (e.g., $a_{1}\rightarrow a_{2}\rightarrow a_{4}$), along with the average problem complexity and text-image semantics of multiple questions sharing this template.
@@ 2.3 Adaptive Reasoning and Verification
During inference, given a multimodal test query $x_{t}$, we compute its PCC and TIS, and perform nearest neighbor matching [35] against pre-constructed thought cards $\mathbb{C}$ to identify the five most relevant cards that best align with its complexity level and semantics. The selection process involves ranking cards based on similarity for each attribute independently:
$R_{\text{TIS}}(c)=\text{Rank}\left(E_{x_{t}}(i_{t},t_{t})^{\top}E_{x}(i_{c},t_{c})\right),\forall c\in\mathbb{C}$ (5)
$R_{\text{PCC}}(c)=\text{Rank}\left(\frac{1}{|PCC_{x_{t}}-PCC_{c}|}\right),\forall c\in\mathbb{C}$ (6)
where $\text{Rank}(\cdot)$ assigns rankings in descending order, with higher values receiving better rankings (e.g., the highest value is ranked 1, the lowest value is ranked $|\mathbb{C}|$). We then combine these rankings to compute a total ranking score:
$R(x_{t},c)=R_{\text{TIS}}(c)+R_{\text{PCC}}(c)$ (7)
Finally, we select the five thought cards with better combined rankings:
$NN_{5}(x_{t},\mathbb{C})=\mathop{\arg\min}_{\mathbb{C}_{x_{t}}\subseteq\mathbb{C},|\mathbb{C}_{x_{t}}|=5}{\textstyle\sum_{c\in\mathbb{C}_{x_{t}}}}R(x_{t},c)$ (8)
where $\mathbb{C}_{x_{t}}\subseteq\mathbb{C}$ contains the five thought cards with the best combined rankings. We instantiate these templates for the test query to obtain five candidate solutions. To identify the best reasoning trajectory, we employ both self-consistency checks and text-domain outcome reward models due to the scarcity of visual-domain verification models. Details are described in Appendix B.4.
In summary, our approach adaptively selects high-level reasoning guidelines based on problem attributes, elegantly integrating the model’s internal implicit reasoning capabilities with external explicit guidance. This dynamic selection mechanism enables more flexible and efficient reasoning across diverse problem types while maintaining robust performance.